Let Xi, X2, X3 be i.id. N(0.1) Suppose Yı = Xi + X2 + X3,Ý, = Xi-X2, у,-X,-X3. Find the joint pdf of Y-(y, Ya, y), using: andom variables. a. The method of variable transformations (Jacobian), b. Multivariate normal distribution properties.
Answer #1

a. The joint distribution of (X1, X2, X3) is f(x1,x2, x3) = Transform (X, X, X3) Vilý%) uhere

then x1= x1 , x2 . x3) Jacobian 1/3 -1皿 皿1-1/3-2/3 | = 3, 皿 1/3 Now

Hence joint distribution of Yi, Y2, Y3 is 3/2 (27)3,2 × 3

where, X2~ N(0, 0-null vector, I3), 13= identity matriz order 3. A = 1-1 3 0 0 A0=0, AA- 0 2 1